# On the Robustness of RSMA to Adversarial BD-RIS-Induced Interference

### v2

@@ I Introduction
RSMA is a split transmission scheme whereby the BS (BS) encodes a part of the users’ data into private messages - transmitted via private precoders - with the remaining users’ data being encoded into a common message - transmitted via a common precoder [10, 11]. On the receiver side, SIC (SIC) is utilized to aid in decoding the individual users’ messages. The private precoders are designed as unicast precoders, which - similarly to SDMA - makes them sensitive to imperfect CSI. The common precoder is constructed as a multicast precoder and, thus, should be more robust to inter-user interference arising due to CSI imperfections. Consequently, by adaptively re-allocating the power between the private and common precoders, RSMA is able to adapt to CSI imperfection [11]. In the conference version of this article [6], we observed that RSMA is able to adapt also in the face of a channel estimation attack induced by a passive RIS, when the CSI available at the BS is imperfect. However, the extent of RSMA’s susceptibility to RIS-induced attacks is still an open challenge with many different parts of the RSMA protocol, including the uplink pilot training, power allocation, and SIC imperfections, requiring deeper investigation. One particularly poignant point is the susceptibility to attacks by novel RIS architectures that feature non-diagonal reflection matrices. These architectures, referred to as BD-RIS, generalize conventional RIS designs by introducing inter-element connections [12]. Such connections enable more advanced optimization capabilities and enhanced energy radiation, which in turn boosts reflection performance at the cost of additional circuit complexity. Nevertheless, the enhanced reflection efficiency of BD-RIS may also strengthen their adversarial capabilities, potentially leading to attacks with increased degradation impact. Identifying the distinct vulnerabilities that BD-RIS may introduce to RSMA systems and the advantages that would motivate an attacker to leverage BD-RIS constitutes a major research gap.
@@
This article reports on our extensive study and analysis of the robustness of the RSMA protocol to attacks against channel acquisition induced by BD-RIS, under different attack modes during the uplink pilot training, varying numbers of reflecting elements, and various levels of CSI and SIC imperfections. In our study, we consider three BD-RIS element network architectures: fully connected, group-connected, and single-connected (corresponding to the diagonal RIS). For each architecture, we propose two different BD-RIS-induced interference attacks that exploit the training protocol employed at the BS, namely random and aligned. In both attacks, during uplink pilot training, the BD-RIS is configured to absorb or reflect with random reflection coefficients. Then, during data transmission, the BD-RIS is configured with newly generated random reflection coefficients (random attack) or with optimized coefficients that maximize the reflected power based on the knowledge of the RIS channels (aligned attack). In each case, the proposed algorithm finds a set of reflection coefficients that satisfy the symmetry and unitary constraints of the BD-RIS.
@@
Our numerical study reveals that, when perfect CSI is available, RSMA behaves similarly to SDMA and thus is highly susceptible to the attack, with BD-RIS group-connected and fully connected architectures inducing larger performance degradations than a single-connected RIS. However, RSMA displays a surprising degree of robustness to the BD-RIS-induced attack when the CSI is imperfect, similarly to what we previously reported for conventional RIS in [6]. When the system is under the BD-RIS-induced attack, the adaptation mechanism of RSMA reacts to the inaccurate CSI by allocating more power to the common message. When the transmit power is low (below 10 dBm) RSMA’s rate is being severely degraded by the attack (on par with SDMA), but as the transmit power increases, RSMA is able to recover from the attack. This trend can be observed across all types of BD-RIS attacks. In order to directly quantify this inherent robustness, we propose the robustness index that we define and discuss in subsequent sections. As we gradually adjust the transmit power, the robustness index for SDMA under attack steadily decreases towards a saturation point. In contrast, RSMA’s robustness index initially declines to a local minimum before beginning a continuous rise, consistently remaining above SDMA’s robustness. This highlights the ability of RSMA to sustain performance under attack, given a sufficiently large transmit power. We also show that the performance degradation induced by group and fully connected architectures becomes increasingly severe compared to a conventional RIS as the number of reflecting elements grows. Our key contributions are as follows:
@@
  • For the aligned interference attack strategy, we develop an algorithm to maximize the unwanted reflected power towards legitimate users. To address the resulting complex optimization problems, we propose to exploit the symmetric structure of the BD-RIS reflection matrices by employing duplication matrices. This approach reformulates the problems into simplified QCQP versions that are efficiently solvable using SVD (SVD), followed by projections onto the unitary manifold via the Takagi factorization.
@@
Notation: Boldface lower-case letters denote vectors and upper-case represent matrices. The $\ell_{2}$ norm of a vector $\mathbf{a}$ is denoted by $\|\mathbf{a}\|_{2}$, and the Frobenius norm of a matrix $\mathbf{A}$ by $\|\mathbf{A}\|_{F}$. The $i$th column of a matrix $\mathbf{A}$ is denoted by $\mathbf{A}]_{:,i}$, the transpose and Hermitian transpose of $\mathbf{A}$ are represented by $\mathbf{A}^{T}$ and $\mathbf{A}^{H}$, respectively, and $\otimes$ represents the Kronecker product. The operator $\mathrm{vec}(\cdot)$ transforms a matrix into a column vector by stacking its columns sequentially, $\mathrm{unvec}(\cdot)$ reshapes a vectorized matrix into its original dimensions, and $\mathrm{vech}(\cdot)$ is the half-vectorization operator, which stacks the elements from the lower triangular part (including the main diagonal) of a square matrix into a column vector. Moreover, $\mathrm{vecd}(\cdot)$ converts the diagonal elements of a square matrix into a column vector, $\mathrm{diag}(\cdot)$ transforms a vector into a diagonal matrix, and $\mathrm{bdiag}(\cdot)$ constructs a block diagonal matrix sequentially formed by the input matrices.
@@ II Related Work
The concept of an adversary that configures RIS to launch attacks against the wireless link was originally proposed in [1]. The proposed attack relied on using the RIS to create cancellation signals. Although the attack has the potential to cause significant power degradation in the receiver, the attack requires highly accurate CSI, which would limit its feasibility. Notably, many works that followed showed that a RIS-induced attack can be made effective with limited or even no CSI [3, 5, 13, 6, 8], as long as it takes advantage of the fact that for many wireless protocols the CSI acquisition is sufficiently separated in time from data transmission. For instance, in [3], it was shown that the channel equalization can be disrupted by the operation of the adversarial RIS that randomly flips its reflection pattern between channel acquisition and symbol transmission. This random flipping is particularly harmful to multiple access methods that rely on multi-antenna systems, such as SDMA. In this case, adversarial RIS increases CSI inaccuracy, making linear precoders ineffective in removing inter-user interference. Exactly this kind of attack with randomly set RIS coefficients, humorously referred to by their authors as the “disco-ball attack”, was proposed in [5, 13]. The optimized version of this attack that aligns the RIS channels to boost interference was proposed and studied in [6, 8]. Other protocols and services, such as beam management [14], physical layer key generation [15, 16], and integrated sensing and communication [17, 18], have also been shown to be vulnerable to the attack with RIS flipping its coefficients between different phases of the wireless protocol. However, many open questions remain in studying RIS-induced attacks, particularly as they relate to identifying protocols and methods that may be robust or resilient to the proposed attacks.
@@
RSMA was introduced as a multiple access strategy that exploits a split transmission mechanism to adaptively treat inter-user interference at either the BS or receiver side, depending on the accuracy of the available CSI [10, 11]. When CSI is perfect, RSMA resorts to linear precoding at the BS that can effectively take care of the inter-user interference. On the other hand, when CSI is inaccurate, RSMA employs multicast transmission and SIC at the receiver side to decode (part of) inter-user interference. There are a variety of benefits that come from this increased flexibility, such as increased spectral and energy efficiency [19], improved outage [20], and reduced latency [21, 22]. RSMA has also been shown to be robust and resilient to user mobility [23], and binary link failures [24]. However, one of the properties gaining the most attention is its inherent robustness and resilience to inaccurate CSI [11], which can also be quantified, following [24]. Since RIS-induced interference attacks inherently rely on manipulating the CSI, the outstanding question is how robust RSMA is to inaccurate CSI that results from adversarial RIS activities.
In the preliminary work described in the conference version of this article, we considered the impact of adversarial diagonal RIS on RSMA [9]. What we observed is the surprising robustness of RSMA to the RIS-induced attack, which manifests itself under imperfect CSI scenarios. In this article, we extend our preliminary study by directly quantifying the robustness of RSMA to adversarial attacks using a new metric that we refer to as the robustness index. We focus our attention on the impact of an attack with a BD-RIS architecture that generalizes conventional (diagonal) RIS. Alternative RIS architectures were considered for adversarial attacks to create non-reciprocal channels [25], and increase their angular span [26]. However, this is the first work that considers BD-RIS and inter-element connections in the context of an adversarial scenario. Consequently, in this paper, we propose a set of new algorithms that find valid reflection coefficients that obey the properties of BD-RIS. Furthermore, we consider different attack modes during the uplink pilot training, and the impact of adaptive power allocation, and SIC imperfections.
@@
III System Model
Consider the multi-user MISO system depicted in Fig. 1, where a BS equipped with a linear array of $M$ antennas communicates with $U$ single-antenna users, represented by the index set $\mathcal{U}=\{1,2,\cdots,U\}$, with the aid of RSMA. The system operates in the presence of an adversarial entity that controls a BD-RIS comprising $D$ reflecting elements. The attacker manipulates the BD-RIS to disrupt the channel estimation phase at the BS, configuring it to either absorb incoming pilots [6, 5] (absorption mode) or reflect them randomly [13] (reflective mode). The impact of the mode choice will be investigated in Section V.
Subsequently, the attacker generates a new set of reflecting coefficients during the data transmission phase, altering the propagation environment in a manner unknown to the BS. This ensures that the BS operates with corrupted CSI, as its estimates from the training phase do not reflect the RIS’s true, active configuration during data transmission. The discrepancy between estimated and actual CSI reduces the performance of precoders, ultimately generating inter-user interference. Fig. 2 illustrates the overall adversarial protocol.
III-A BD-RIS Architectures
We consider two BD-RIS types, the fully connected and group-connected architectures. In the fully connected RIS, all elements are interconnected through reconfigurable impedances. A fully connected RIS with $D$ reflecting elements operates as a $D$-port reciprocal network. The interaction among elements is captured by the scattering matrix $\bm{\Theta}\in\mathbb{C}^{D\times D}$, which characterizes the reflection behavior of the system. Energy conservation imposes the constraint $\bm{\Theta}\bm{\Theta}^{H}\preceq\mathbf{I}_{D}$, implying that the Frobenius norm satisfies $\frac{1}{\sqrt{D}}|\bm{\Theta}|_{F}\leq 1$. When the circuit is lossless, the scattering matrix becomes unitary, i.e., $\bm{\Theta}\bm{\Theta}^{H}=\mathbf{I}_{D}$. Furthermore, due to the reciprocal nature of the BD-RIS hardware, the symmetry condition $\bm{\Theta}=\bm{\Theta}^{T}$ must also be satisfied. This property implies that $\frac{D(D+1)}{2}$ independent coefficients are needed to fully define the entries of $\bm{\Theta}$.
@@ III-A BD-RIS Architectures
where each submatrix $\bm{\Theta}_{g}\in\mathbb{C}^{D_{g}\times D_{g}}$ models the signal reflections induced by its corresponding fully connected group, with $D_{g}$ denoting the number of elements of the $g$th group, such that $D=\sum_{g}^{G}D_{g}$. Similar to the fully connected case, each submatrix satisfies the conditions $\bm{\Theta}_{g}=\bm{\Theta}_{g}^{T}$ and $\bm{\Theta}_{g}\bm{\Theta}_{g}^{H}\preceq\mathbf{I}_{D_{g}}$. The number of independent parameters needed to configure each group is, therefore, $\frac{D_{g}(D_{g}+1)}{2}$, in which $D_{g}\leq D$. From this point onward, we consider lossless reflection models, i.e., we assume $\bm{\Theta}\bm{\Theta}^{H}=\mathbf{I}_{D}$ and $\bm{\Theta}_{g}\bm{\Theta}_{g}^{H}=\mathbf{I}_{D_{g}}$ for the fully and group-connected RIS, respectively. The impact of different adversarial strategies is analyzed in the following sections.

### v3

@@ I Introduction
RSMA is a split transmission scheme whereby the BS (BS) encodes a part of the users’ data into private messages - transmitted via private precoders - with the remaining users’ data being encoded into a common message - transmitted via a common precoder [10, 4]. On the receiver side, SIC (SIC) is utilized to aid in decoding the individual users’ messages. The private precoders are designed as unicast precoders, which - similarly to SDMA - makes them sensitive to imperfect CSI. The common precoder is constructed as a multicast precoder and, thus, should be more robust to inter-user interference arising due to CSI imperfections. Consequently, by adaptively re-allocating the power between the private and common precoders, RSMA is able to adapt to CSI imperfection [4]. In the conference version of this article [7], we observed that RSMA is able to adapt also in the face of a channel estimation attack induced by a passive RIS, when the CSI available at the BS is imperfect. However, the extent of RSMA’s susceptibility to RIS-induced attacks is still an open challenge with many different parts of the RSMA protocol, including the uplink pilot training, power allocation, and SIC imperfections, requiring deeper investigation. One particularly poignant point is the susceptibility to attacks by novel RIS architectures that feature non-diagonal reflection matrices. These architectures, referred to as BD-RIS, generalize conventional RIS designs (i.e., diagonal RIS) by introducing inter-element connections [33]. Such connections enable more advanced optimization capabilities and enhanced energy radiation, which in turn boosts reflection performance at the cost of additional circuit complexity. Nevertheless, the enhanced reflection efficiency of BD-RIS may also strengthen their adversarial capabilities, potentially leading to attacks with increased degradation impact. Identifying the distinct vulnerabilities in RSMA systems and the advantages that would motivate an attacker to leverage BD-RIS constitutes a major research gap.
@@
This article reports on our extensive study and analysis of the robustness of the RSMA protocol to attacks against channel acquisition induced by BD-RIS, under different attack modes during the uplink pilot training, varying numbers of reflecting elements, and various levels of CSI and SIC imperfections. Unlike diagonal RIS attacks, where the adversary is limited to independent per-element phase shifts, BD-RIS-based attacks exploit inter-element coupling through non-diagonal scattering matrices, causing stronger perturbations to channel estimates than a conventional RIS of the same dimension (or, equivalently, achieving similar degradation with fewer reflecting elements). In our study, we consider three BD-RIS element network architectures: fully connected, group-connected, and single-connected (corresponding to the diagonal RIS). For each architecture, we propose two different BD-RIS-induced interference attacks that exploit the training protocol employed at the BS, namely random and aligned. In both attacks, during uplink pilot training, the BD-RIS is configured to absorb or reflect with random reflection coefficients. Then, during data transmission, the BD-RIS is configured with newly generated random reflection coefficients (random attack) or with optimized coefficients that maximize the reflected power based on the knowledge of the RIS channels (aligned attack). In each case, the proposed algorithm finds a set of reflection coefficients that satisfy the symmetry and unitary constraints of the BD-RIS.
@@
Our numerical study reveals that, when perfect CSI is available, RSMA behaves similarly to SDMA and thus is highly susceptible to the attack (corroborating the findings reported in [21, 8, 46, 41]), with BD-RIS group-connected and fully connected architectures inducing larger performance degradations than a single-connected RIS. However, even though the BS is assumed to be completely unaware of the BD-RIS-induced manipulation, RSMA displays a surprising degree of robustness when the CSI is imperfect, similarly to what we previously reported for conventional RIS in [7]. The adaptation mechanism of RSMA reacts to the resulting inaccurate CSI by allocating more power to the common message and, as a consequence, unintentionally mitigates the impact of the BD-RIS-induced attack. When the transmit power is low (below 10 dBm), RSMA’s rate is severely degraded by the attack (on par with SDMA), but as the transmit power increases, RSMA is able to recover from the attack. This trend can be observed across all types of BD-RIS attacks. In order to directly quantify this inherent robustness, we propose the robustness index, which we define and discuss in subsequent sections. As we gradually adjust the transmit power, the robustness index for SDMA under attack steadily decreases towards a saturation point. In contrast, RSMA’s robustness index initially declines to a local minimum before beginning a continuous rise, consistently remaining above SDMA’s robustness. This highlights the ability of RSMA to sustain performance under attack, given a sufficiently large transmit power. We also show that the performance degradation induced by group and fully connected architectures becomes increasingly severe compared to a conventional RIS as the number of reflecting elements grows [^1] This work shows that BD-RIS-induced attacks can cause significant performance degradation. However, it is noteworthy that practical BD-RIS implementations may also be affected by hardware impairments, including phase shift errors (e.g., [35]), mutual coupling (e.g., [52]), and non-ideal response matrix (e.g., [5]). These impairments may also influence the effectiveness of adversarial attacks in practice. A comprehensive investigation of the impact of hardware impairments on adversarial effectiveness is left for future work.. Our key contributions are as follows:
@@
  • For the aligned interference attack strategy, we develop an algorithm to maximize the unwanted reflected power towards legitimate users. To address the resulting complex optimization problem, we propose exploiting the symmetric structure of the BD-RIS reflection matrices by using duplication matrices. This approach reformulates the problem into a simplified QCQP that can be efficiently solved using SVD (SVD), followed by projections onto the unitary manifold via the Takagi factorization.
@@
Notation: Boldface lower-case letters denote vectors and upper-case letters represent matrices. The $\ell_{2}$ norm of a vector $\mathbf{a}$ is denoted by $\|\mathbf{a}\|_{2}$, and the Frobenius norm of a matrix $\mathbf{A}$ by $\|\mathbf{A}\|_{F}$. The $i$th column of a matrix $\mathbf{A}$ is denoted by $\mathbf{A}]_{:,i}$, the transpose and Hermitian transpose of $\mathbf{A}$ are represented by $\mathbf{A}^{T}$ and $\mathbf{A}^{H}$, respectively, and $\otimes$ represents the Kronecker product. The operator $\mathrm{vec}(\cdot)$ transforms a matrix into a column vector by stacking its columns sequentially, $\mathrm{unvec}(\cdot)$ reshapes a vectorized matrix into its original dimensions, and $\mathrm{vech}(\cdot)$ is the half-vectorization operator, which stacks the elements from the lower triangular part (including the main diagonal) of a square matrix into a column vector. Moreover, $\mathrm{vecd}(\cdot)$ converts the diagonal elements of a square matrix into a column vector, $\mathrm{diag}(\cdot)$ transforms a vector into a diagonal matrix, and $\mathrm{bdiag}(\cdot)$ constructs a block diagonal matrix sequentially formed by the input matrices.
@@ II Related Work
The concept of an adversary that configures RIS to launch attacks against the wireless link was originally proposed in [37]. The proposed attack relied on using the RIS to create cancellation signals. Although the attack has the potential to cause significant power degradation at the receiver, it requires highly accurate CSI, which would limit its feasibility. Notably, many works that followed showed that a RIS-induced attack can be made effective with limited or even no CSI [45, 21, 18, 8, 41], as long as it takes advantage of the fact that for many wireless protocols the CSI acquisition is sufficiently separated in time from data transmission. For instance, in [45], it was shown that the channel equalization can be disrupted by the operation of the adversarial RIS that randomly flips its reflection pattern between channel acquisition and symbol transmission. This random flipping is particularly harmful to multiple access methods that rely on multi-antenna systems, such as SDMA. In this case, adversarial RIS increases CSI inaccuracy, making linear precoders ineffective in removing inter-user interference. Exactly this kind of attack with randomly set RIS coefficients, humorously referred to by their authors as the “disco-ball attack”, was proposed in [21, 18]. The optimized version of this attack, which aligns the RIS channels to boost interference, was proposed and studied in [8, 41]. Other protocols and services, such as beam management [16], physical layer key generation [29, 48], and integrated sensing and communication [19, 42], have also been shown to be vulnerable to the attack with RIS flipping its coefficients between different phases of the wireless protocol. However, many open questions remain in studying RIS-induced attacks, particularly as they relate to identifying protocols and methods that may be robust or resilient to the proposed attacks.
@@
RSMA was introduced as a multiple access strategy that exploits a split transmission mechanism to adaptively treat inter-user interference at either the BS or receiver side, depending on the accuracy of the available CSI [10, 4]. When CSI is perfect, RSMA resorts to linear precoding at the BS that can effectively take care of the inter-user interference. On the other hand, when CSI is inaccurate, RSMA employs multicast transmission and SIC at the receiver side to decode (part of) inter-user interference. There are a variety of benefits that come from this increased flexibility, such as increased spectral and energy efficiency [53], improved outage [36], and reduced latency [50, 39]. RSMA has also been shown to be robust and resilient to user mobility [13], binary link failures [40], and, as hinted at earlier, inaccurate CSI [4].
The benign application of RIS to RSMA has been extensively investigated in the literature. In [10], the interplay between RSMA and RIS was discussed, highlighting how rate splitting can exploit the additional degrees of freedom introduced by programmable propagation environments. Several works have also considered joint design and optimization of RSMA and RIS. For instance, [25] studied uplink RSMA in multi-RIS-aided millimeter-wave systems with joint optimization of user clustering, power allocation, and active and passive beamforming. The work in [11] investigated RSMA assisted by simultaneous transmitting and reflecting RIS over spatially correlated channels. More recently, [39] considered the joint optimization of transmit precoding, power allocation, and RIS coefficients under finite blocklength assumptions, while [26] considered RIS-assisted RSMA in integrated sensing and communication networks. All these works use RIS as a performance-enhancing technology and focus on joint transmission design under channel uncertainty models arising from conventional channel estimation and acquisition processes.
In adversarial settings, the RIS acts as an illegitimate component beyond the control of the BS, preventing the joint RIS-RSMA transmission design. It is thus important to consider how RSMA, which explicitly adapts its interference management strategy to channel conditions, behaves when the propagation environment is deliberately manipulated by an illegitimate RIS. In a preliminary conference study, we were the first to investigate this question by examining the behavior of RSMA in the presence of an adversarial single-connected RIS [7]. There, we observed that RSMA remains surprisingly robust to RIS-induced attacks when already operating with imperfect channel knowledge. This naturally motivates extending the analysis to more general and powerful BD-RIS architectures.
On the BD-RIS front, despite being a relatively recent concept, it has attracted significant attention from the research community. The foundational work in [43] modeled RIS as a multi-port network using scattering-parameter analysis, enabling the characterization of inter-element coupling and structural constraints beyond diagonal architectures. Building on this perspective, [31] investigated single-connected, group-connected, and fully connected BD-RIS architectures, while a broader overview of their role in future wireless systems was presented in [33]. Reduced-complexity non-diagonal RIS designs that approach the performance of fully connected architectures were further studied in [34], and low-complexity beamforming methods for BD-RIS-aided multi-user networks were proposed in [15]. The integration of BD-RIS with RSMA has also been investigated. In [14], fully connected RIS architectures were considered and jointly optimized with RSMA transmit beamformers to maximize the downlink sum rate. Robust BD-RIS-assisted RSMA designs under imperfect CSI were studied in [32], while [44] investigated BD-RIS-assisted RSMA under finite blocklength constraints for URLLC (URLLC) scenarios. Uplink RSMA with BD-RIS was recently considered in [27] using learning-based optimization, and BD-RIS-assisted RSMA has also been studied in the context of simultaneous wireless information and power transfer in [3].
BD-RIS architectures have also been investigated from an adversarial and security-oriented perspective. In [47] and [48], non-diagonal RIS were shown to enable channel reciprocity attacks and to disrupt physical-layer key generation without active jamming. Passive omnidirectional jamming using “DISCO surfaces” was proposed in [20]. More recently, [22] analyzed worst-case adversarial channels induced by BD-RIS and identified fundamental limits of simplified surface architectures. Although these works establish the adversarial potential of BD-RIS, the impact of such attacks on next-generation multiple access schemes, particularly on the robustness of RSMA, remains largely unexplored.
In the conference version of this work, we took a first step toward addressing this gap by studying the impact of adversarial diagonal RIS on RSMA [7]. In this article, we extend our preliminary study by quantifying the robustness of RSMA to adversarial BD-RIS attacks. The details of the RSMA system model, BD-RIS architectures, and the considered attack mechanisms are introduced next.
III System Model
Consider the multi-user MISO system depicted in Fig. 1, where a BS equipped with a linear array of $M$ antennas communicates with $U$ single-antenna users, represented by the index set $\mathcal{U}=\{1,2,\ldots,U\}$, with the aid of RSMA. The system operates in the presence of an adversarial entity that controls a BD-RIS comprising $D$ reflecting elements. The attacker manipulates the BD-RIS to disrupt the channel estimation phase at the BS, configuring it to either absorb incoming pilots [8, 21] (absorption mode) or reflect them randomly [18] (reflective mode). The impact of the mode choice will be investigated in Section V.
Subsequently, the attacker generates a new set of reflecting coefficients during the data transmission phase, altering the propagation environment in a manner unknown to the BS. This ensures that the BS operates with corrupted CSI, as its estimates from the training phase do not reflect the RIS’s true configuration during data transmission. The discrepancy between estimated and actual CSI reduces the performance of precoders, ultimately leading to inter-user interference. Fig. 3 illustrates the overall adversarial protocol.
Figure 2: Illustrative comparison of different RIS architectures with four elements: (a) single-connected; (b) group-connected; and (c) fully connected. $Z_{i}$ represents the reconfigurable self-impedance of element $i$, while $Z_{ij}$ denotes the mutual impedance between elements $i$ and $j$, with $i,j\in\{1,\ldots,4\}$.
III-A BD-RIS Architectures
We consider two BD-RIS architectures: fully connected and group-connected. Fig. 2 illustrates the structural differences between these BD-RIS architectures and the conventional (single-connected) RIS. Specifically, in the fully connected RIS, all elements are interconnected through reconfigurable impedances. A fully connected RIS with $D$ reflecting elements operates as a $D$-port reciprocal network. The interaction among elements is captured by the scattering matrix $\bm{\Theta}\in\mathbb{C}^{D\times D}$, which characterizes the reflection behavior of the system. Energy conservation imposes the constraint $\bm{\Theta}\bm{\Theta}^{H}\preceq\mathbf{I}_{D}$, implying that the Frobenius norm satisfies $\frac{1}{\sqrt{D}}|\bm{\Theta}|_{F}\leq 1$. When the circuit is lossless, the scattering matrix becomes unitary, i.e., $\bm{\Theta}\bm{\Theta}^{H}=\mathbf{I}_{D}$. Furthermore, due to the reciprocal nature of the BD-RIS hardware, the symmetry condition $\bm{\Theta}=\bm{\Theta}^{T}$ must also be satisfied [^2] The concept of non-reciprocal BD-RIS, featuring non-symmetric unitary reflection matrices, has also been proposed recently [30]. Such architectures can break channel reciprocity and induce a mismatch between uplink and downlink cascaded channels even if the RIS configuration is kept fixed over time, thereby facilitating adversarial attacks.. This property implies that $\frac{D(D+1)}{2}$ independent coefficients are needed to fully define the entries of $\bm{\Theta}$.
@@ III-A BD-RIS Architectures
where each submatrix $\bm{\Theta}_{g}\in\mathbb{C}^{D_{g}\times D_{g}}$ models the signal reflections induced by its corresponding fully connected group, with $D_{g}$ denoting the number of elements of the $g$th group, such that $D=\sum_{g}^{G}D_{g}$. Similar to the fully connected case, each submatrix satisfies the conditions $\bm{\Theta}_{g}=\bm{\Theta}_{g}^{T}$ and $\bm{\Theta}_{g}\bm{\Theta}_{g}^{H}\preceq\mathbf{I}_{D_{g}}$. The number of independent parameters needed to configure each group is, therefore, $\frac{D_{g}(D_{g}+1)}{2}$, in which $D_{g}\leq D$. From this point onward, we consider ideal reflection models, i.e., we assume $\bm{\Theta}\bm{\Theta}^{H}=\mathbf{I}_{D}$ and $\bm{\Theta}_{g}\bm{\Theta}_{g}^{H}=\mathbf{I}_{D_{g}}$ for the fully and group-connected RIS, respectively. The impact of different adversarial strategies is analyzed in the following sections.